OpenStudy (anonymous):

Please find the area enclosed by this parametric function: \[\begin{gathered} x = 2a\cos t - a\cos 2t \hfill \\ y = 2a\sin t - a\sin 2t \hfill \\ \end{gathered} \]

OpenStudy (amistre64):

the area of a section of a circle is 1/2 pi t^2 right?

OpenStudy (amistre64):

2pi --- r^2 .. yeah 2

OpenStudy (anonymous):

right

OpenStudy (amistre64):

and r here is the function provided

OpenStudy (anonymous):

Could you show me the details?

OpenStudy (amistre64):

if i could remember the details i would :)

OpenStudy (anonymous):

But I don't think your answer is right...

OpenStudy (amistre64):



OpenStudy (amistre64):

my answer is neither right nor wrong, its simply incomplete

OpenStudy (amistre64):

im thinking this is a double integral

OpenStudy (amistre64):

dr dt

OpenStudy (amistre64):

we measure the value or r thru each degree, and add up all the degrees and their little rs

OpenStudy (amistre64):

have you done doubles yet?

OpenStudy (amistre64):

r^2 = x^2 + y^2 so that part is "simple" enough ...

OpenStudy (anonymous):

Yes I have learn

OpenStudy (amistre64):

\[r^2 = (2a\ cos(t)−acos(2t))^2\ +(2a\ sin(t)−a\ sin(2t))^2\] unless im reading the material wrong

OpenStudy (amistre64):

t = 0 to 2pi

OpenStudy (amistre64):

i assume "a" is considered to be a constant?

OpenStudy (anonymous):

\[dA = \frac{1} {2}r*rdt\] Is it you are thinking?

OpenStudy (amistre64):

yes, or something vaguely similar to it..